OpenStudy (anonymous):

Question 9 (Multiple Choice Worth 1 points) [5.06] MA.7.A.1.2 Fayed tipped the pizza delivery driver three dollars for pizzas that cost eighteen dollars. Approximately what percent did he tip the driver? Round your answer to the nearest percent. three percent six percent fifteen percent seventeen percent Question 10 (Multiple Choice Worth 1 points) [5.01] MA.6.A.2.2 Four friends went scuba diving. According to the following statements, who spotted sea life at the fastest rate? Lance saw 2 reef sharks in 6 minutes. Jasmine saw 3 sea cucumbers in 8 minutes. Ro

OpenStudy (anonymous):

18/100= 1 percent 3/1 percent= your answer

OpenStudy (anonymous):

8/3=rate for Jasmine 6/2=rate for Lance pick the highest for your answer..

OpenStudy (anonymous):

@aqua666 what do u mean ??